OpenStudy (anonymous):

Solve radical expression 5x+√ x+7-13=0

OpenStudy (anonymous):

First you have to isolate the radical on a side. So subtract 5x from both sides and also add 13 to both sides. Then you square both sides. (this will require you to FOIL on the side with the 13-5x). The resulting equation will be quadratic, so you can solve that by getting all terms to one side and then trying to factor it or using hte quadratic formula. lastly, you HAVE to check your answers because its possible to get answers that don't really work here.

OpenStudy (anonymous):

\[\sqrt{(x+7)}=13-5x\] \[(\sqrt{(x+7)})^{2}=(13-5x)^{2}\] \[x+7=(13-5x)(13-5x)\] etc.

OpenStudy (anonymous):

x+7=25x^2-130x+169

OpenStudy (anonymous):

\[x+7=169-65x-65x+25x ^{2}\]

OpenStudy (anonymous):

\[0=25x ^{2}-130x+169-x-7\]

OpenStudy (anonymous):

\[0=25x ^{2}-131x+162\]

OpenStudy (anonymous):

Wow - those are big numbers....I'd use the quadratic formula.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

x=2